OpenStudy (gabyramirezzz):

For the function y= -1+6cos(2pi/7(x-5)) what is the maximum value? (sorry I keep asking so many questions)

jimthompson5910 (jim_thompson5910):

cosine is a lot like sine in many ways (hence the name coSINE) cosine has a range from -1 to 1, ie, \[\Large -1 \le \cos(x) \le 1\]

jimthompson5910 (jim_thompson5910):

use the same idea used last time and hopefully you see how to find the range of the entire function

OpenStudy (gabyramirezzz):

oh, so -6<cos<6?

jimthompson5910 (jim_thompson5910):

more like \[\Large -6 \le 6\cos(x) \le 6\]

OpenStudy (gabyramirezzz):

oh

jimthompson5910 (jim_thompson5910):

therefore the max of the piece "6cos(2pi/7(x-5))" is 6 that is the largest output you can get from that piece

jimthompson5910 (jim_thompson5910):

so what is the overall max?

OpenStudy (gabyramirezzz):

6? or so we have to do something with the -1?

jimthompson5910 (jim_thompson5910):

yeah that -1 will change things

OpenStudy (gabyramirezzz):

so is it -6?

jimthompson5910 (jim_thompson5910):

-1 + 6 = 5 actually

jimthompson5910 (jim_thompson5910):

the -1 pulls the graph down 1 unit

jimthompson5910 (jim_thompson5910):

taking the max down 1 unit as well

OpenStudy (gabyramirezzz):

ohh

OpenStudy (gabyramirezzz):

so the answer is five?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (gabyramirezzz):

thank you again!!